### v1

# Visual-Friendly Concept Protection via Selective Adversarial Perturbations

Abstract.
Personalized concept generation by tuning diffusion models with a few images raises potential legal and ethical concerns regarding privacy and intellectual property rights. Researchers attempt to prevent malicious personalization using adversarial perturbations. However, previous efforts have mainly focused on the effectiveness of protection while neglecting the visibility of perturbations. They utilize global adversarial perturbations, which introduce noticeable alterations to original images and significantly degrade visual quality. In this work, we propose the Visual-Friendly Concept Protection (VCPro) framework, which prioritizes the protection of key concepts chosen by the image owner through adversarial perturbations with lower perceptibility. To ensure these perturbations are as inconspicuous as possible, we introduce a relaxed optimization objective to identify the least perceptible yet effective adversarial perturbations, solved using the Lagrangian multiplier method. Qualitative and quantitative experiments validate that VCPro achieves a better trade-off between the visibility of perturbations and protection effectiveness, effectively prioritizing the protection of target concepts in images with less perceptible perturbations.
Unauthorized personalizations, adversarial examples, diffusions, privacy protections, copyright protections
copyright: acmlicensedjournal: TOMMdoi: XXXXXXX.XXXXXXXccs: Information systems Multimedia information systemsccs: Security and privacy Human and societal aspects of security and privacy
1. Introduction
With the advent of popular generative models (Ho et al., 2020; Rombach et al., 2022; Song et al., 2020) such as Stable Diffusion (Rombach et al., 2022) and DALL·E 2/3 (Ramesh et al., 2022), people lacking expertise in drawing or photography can effortlessly create realistic or artistic works using simple textual descriptions. However, the success of these models has raised significant concerns about privacy, intellectual property rights, and various legal and ethical issues. For example, an adversary could easily generate sensitive specific concepts, such as personal fake images, or imitate renowned artworks for commercial purposes, using only a few reference images and concept personalization techniques such as Textual Inversion (Gal et al., 2022) or DreamBooth (Ruiz et al., 2023).
Figure 1. The first row presents protected images generated by Anti-DreamBooth (Le et al., 2023), Mist (Zhang et al., 2022), MetaCloak (Liu et al., 2024), SDS(-) (Xue et al., 2024), PhotoGuard (Salman et al., 2023), and our method VCPro. The second row shows magnification sections of the background and face regions from different protected images. The third row displays the result of applying Textual Inversion to the protected data. PSNR and SSIM evaluate the visibility of adversarial perturbations; higher values indicate better results. For a fair comparison, all adversarial perturbations have a size of 8/255. Our method achieves a superior balance between final protection effectiveness and the visual quality of perturbations, prioritizing the protection of user-important concepts compared to previous works.
One prevailing direction to mitigate these potential risks is to leverage adversarial attack techniques, transforming original images into adversarial examples, termed “protected images”. These protected images can misguide the personalized diffusion model and deceive its generation process, which resists malicious editing or personalization (Salman et al., 2023; Liang et al., 2023; Liang and Wu, 2023; Le et al., 2023; Shan et al., 2023; Liu et al., 2024). For instance, AdvDM (Liang et al., 2023) employs adversarial attacks in the inversion stage of Stable Diffusion to safeguard against malicious imitation of the style of a specific artist. Later, its updated version, Mist (Liang and Wu, 2023), enhances the protection efficacy of protected images by adding a textual loss, extending its application from Textual Inversion to DreamBooth. Unlike AdvDM and Mist, which focus on art style protection, Anti-DreamBooth (Le et al., 2023) undermines DreamBooth model generation quality to enhance privacy by adding adversarial perturbations to human face images before posting online. Furthermore, MetaCloak (Liu et al., 2024) leverages a meta-learning strategy and data transformations to generate more effective and robust protected images against DreamBooth.
However, these methods primarily focus on preventing the personalization methods from generating high-quality corresponding images, often resulting in noticeable and unacceptable perturbations. They usually use $11/255$ (Le et al., 2023; Shan et al., 2023; Liu et al., 2024) or $17/255$ (Liang and Wu, 2023) as perturbation size in their paper, which is generally unacceptable for owners of face photos, significantly limiting the usability of these methods in real-world applications. Therefore, we pose the following question in this work: How can we find a better trade-off between visibility of perturbations and protection effectiveness?
To answer this question, we highlight the sparsity of images under concept protection tasks: The critical information that deserves protection constitutes only a part of the image. Previous approaches aim to protect the entire image, including background regions and other non-essential information, enhancing the visibility of perturbations. As shown in Fig. 1, the protected images generated by Mist, Anti-DreamBooth, MetaCloak, SDS(-), and PhotoGuard, exhibit noticeable odd textures on the entire image: face, neck, and background, and their final protective effects akin to a special style picture of the target person. In contrast, we prioritize protecting essential information within an image, utilizing a more stealthy adversarial perturbation.
To this end, we introduce a Visual-Friendly Concept Protection (VCPro) framework to counteract unauthorized concept-driven text-to-image synthesis. This framework learns selective adversarial perturbations targeting important regions. Unlike discriminative tasks where important information is class-related and provided by target model gradients, identifying crucial information in generative tasks is challenging. VCPro utilizes user-provided masks for target concept protection, which can be supplemented by other privacy detection tools for online platforms. In that way, we propose a regional adversarial loss using spatial information to focus on selected areas. To further enhance visual quality, we apply a Lagrangian multiplier-based solution, shifting from maximizing protection effectiveness to minimizing perceptibility while ensuring effective protection. Considering human sensitivity to low-frequency changes, we measure perturbation perceptibility in the frequency domain. Experiments on models like Textual Inversion and DreamBooth validate VCPro’s effectiveness. Our approach yields subtler adversarial perturbations compared to baselines like Mist and Anti-DreamBooth, especially FID, which is reduced from $96.24$ to $27.03$.
Our contributions are summarized as follows:
  • We point out that the existing image protection methods over-emphasize the final protection effectiveness while neglecting the visual appearance of the protected images.
  • We propose a visual-friendly concept protection framework to prioritize protecting essential information within images by a regional adversarial loss and loosen the optimization objective from finding the strongest perturbations within the perturbation size to the smallest feasible perturbations.
  • Experiments demonstrate that our approach can focus on crucial concepts specified by users with lower perceptibility than baselines, achieving a better trade-off between protection effectiveness and perturbation visibility.
2. Related Work
Personalization of Diffusions Models. Personalization for specific concepts (attributions, styles, or objects) has been a long-standing goal in the image generation field. In text-to-image diffusion models, previous researchers have primarily concentrated on prompt learning and test-time tuning of pre-trained models to generate images based on specific target concepts using special language tokens. Textual Inversion adjusts text embeddings of a new pseudoword to describe the concept (Gal et al., 2022). DreamBooth fine-tunes denoising networks to connect the novel concept and a less commonly used word token (Ruiz et al., 2023). Based on that, more works (Voynov et al., 2023; Zhang et al., 2023a; Kumari et al., 2023) are proposed to improve controllability and flexibility in processing image visual concepts. In this paper, we have selected Textual Inversion and DreamBooth as the techniques used by the adversary due to their popularity and representativeness.
Imperceptibility Adversarial Attack. Adversarial examples (Szegedy et al., 2013; Carlini and Wagner, 2017; Duan et al., 2021; Zhang et al., 2022; Mi et al., 2023) are initially introduced by adding imperceptible noise to original data, fooling classifiers into misclassifying with high confidence. Recently, more and more researchers aim to improve the imperceptibility of adversarial examples, and they make use of a variety of tools such as perceptual color distance (Zhao et al., 2021), low-frequency spatial constraints (Luo et al., 2022), hybrid attacks in frequency and spatial domain (Jia et al., 2022), and invertible neural networks (Chen et al., 2023), etc. But they are mainly aimed at discriminative tasks such as image classification, where important information in the image can be fed back relatively accurately by the gradient of the target model, whereas we target a diffusion-based generative model whose gradient is also for the whole image.
Adversarial Examples Against Unauthorized AI Generation. Unauthorized AI Generation seriously harms the safety of society and researchers are looking at several angles to reduce its risk (Zhang et al., 2024; Becattini et al., 2023; Mahto et al., 2023). Adversarial attacks are a promising way against unauthorized AI generation (Ruiz et al., 2020; Wang et al., 2022a, b; Zhu et al., 2023; Tang et al., 2024). Ruiz et al. (2020) are the first to adapt adversarial attack methods to undermine the generation of realistic deepfake images, thereby enhancing privacy security. Anti-Forgery (Wang et al., 2022a) computes adversarial noise in the Lab color space to enhance the imperceptibility and robustness of various image transformations. Unlike others, DeepFake Disrupter (Wang et al., 2022b) generates adversarial noise using an autoencoder instead of gradients and spoofs deepfake generators while remaining undetected by deepfake detectors. Zhu et al. (2023) utilize adversarial examples to protect images while simultaneously recovering original image information from tampered images. In contrast to discriminative tasks, adversarial examples play a positive role in the image generation field.
With the rising popularity of diffusion models, previous works are difficult to adapt due to the obvious differences between diffusion and GAN or other generation models. Therefore, there is a growing interest in using adversarial examples to protect images from being maliciously exploited by diffusion models (Salman et al., 2023; Liang et al., 2023; Liu et al., 2024). Photoguard (Salman et al., 2023) attacks the VAE encoder or whole model to prevent images from being maliciously edited by diffusion models in the test time. At the same time, AdvDM (Liang et al., 2023) employs adversarial examples to protect artists from having their styles imitated by textual inversion by maximizing the denoising loss of a pre-trained diffusion model. As an upgraded version of AdvDM, Mist (Liang and Wu, 2023) adds a texture loss that measures the distance between the encoder representation of the original image and the perturbed image in the image encoder of latent diffusion. Similar to them, GLAZE (Shan et al., 2023) proposes to measure the art style similarity between the perturbed image and the target style using a pre-trained style-transfer model, and adversarial examples can protect the art style-related attributes in the original image directly. Different from them, Anti-DreamBooth (Le et al., 2023) protects privacy information from being learned by DreamBooth. Based on that, MetaCloak (Liu et al., 2024) uses a meta-learning strategy and data transformations to generate better effective and robust adversarial examples. Meanwhile, Xue et al. (2024) introduce Score Distillation Sampling (SDS) loss into the perturbation optimization process to reduce computation cost and find maximize or minimize SDS loss both achieve competitive protection effects.
However, except for GLAZE, existing work protects all information for the whole image, which often requires significant noise, significantly affects the user experience, and may overemphasize the background, omitting the information that needs to be protected. In our work, we propose to prioritize limited noise to protect important information in the image, such as faces, and specific IPs. Unlike GLAZE which targets art style, we target specific semantic regions.
3. Preliminaries
Personalization based on Diffusion. Concept-driven personalization customizes the outputs of a generative model to align with a specific concept, such as attributes, style, objects, etc. Most personalization techniques are applied to the latent diffusion model (LDM, parameterized by $\theta$). A typical LDM consists of an image encoder $\mathcal{E}_{\theta}$, an image decoder $\mathcal{D}_{\theta}$, a condition encoder $\tau_{\theta}$, and a conditional denoising UNet $\varepsilon_{\theta}$. The training of LDMs contains diffusion and denoising processes, which both happen in the latent space. For an image $x$, $z_{0}=\mathcal{E}_{\theta}(x)$ is the latent code of $x$. The diffusion process randomly samples time $t$ and Gaussian noise $g$, and latent vector $z_{0}$ is added noise $g$ to $z_{t}$. Then the denoising UNet $\varepsilon_{\theta}$ is assigned to remove the noise added to the latent code $z_{t}$ conditioned by $\tau_{\theta}(y)$ in each timestep $t$, and we set the output of $\varepsilon_{\theta}$ as $g^{\prime}$. The training objective of this LDM is:
(1) $\mathcal{L_{\theta}}:=\mathbb{E}_{z\sim\mathcal{E}(x),y,g\sim\mathcal{N}(0,1),t}\left[\left\|{g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y)\right)}\right\|_{2}^{2}\right],$
where $t$ is a timestep of the diffusion process, and $y$ is the given condition for generation.
Based on LDM, Textual Inversion, and DreamBooth are two main techniques for personalization. Textual Inversion learns new special tokens containing specific concepts to store new knowledge from images. In Textual Inversion, a pseudo token $sks$ and its corresponding embedding vector $v\in\mathcal{P}$ are introduced to describe the concept, where $\mathcal{P}$ is the word embedding space of the text encoder. They optimize $v$ directly by minimizing the objective specified in Eq. (1)s using a small set of images representing the concept as training data.
(2) $\arg\min_{v}\mathbb{E}_{z,y,g,t}\left[\left\|g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y,v)\right)\right\|_{2}^{2}\right].$
DreamBooth also binds new conceptual knowledge to specific words. Unlike Textual Inversion, it fine-tunes part of the parameters of LDMs on training images and class examples $x^{p}$, which is generated from original LDM with prior prompt $y^{p}$ to mitigate catastrophic forgetting.
(3) $\begin{gathered}\theta:=\arg\min_{\theta}\mathbb{E}_{z,y,g,t}\Bigg{[}\left\|g-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y,v)\right)\right\|_{2}^{2}+\left\|g-\varepsilon_{\theta}\left(z^{p}_{t},t^{p},\tau_{\theta}(y^{p})\right)\right\|_{2}^{2}\Bigg{]}.\end{gathered}$
Protected image for Diffusions. Protected images (adversarial examples) aim to add an imperceptible adversarial perturbation $\delta$ to the original image $x$ to make the full process of condition generations ineffective. The protected image $\mathcal{L_{\theta}}$ can be formalized as:
(4) $\begin{gathered}x^{\prime}:=arg\max_{x^{\prime}}\mathcal{L_{\theta}}(x^{\prime},y),\\ \text{ s.t. }\quad\left\|x-x^{\prime}\right\|\leq\epsilon,\end{gathered}$
where $\epsilon$ is the perturbation size, also known as the perturbation budget, and limits the maximum range of adversarial perturbation. Adversarial attack transforms original images into special samples that are difficult to denoise. They enhance the optimization challenge of LDMs, thereby achieving image protection.
4. Visual-Friendly Concept Protection
4.1. Overview
Fig. 2 shows the pipeline of the proposed framework for visual-friendly concept protection. Accurately describing spatial positions through language can be challenging for users, but it can be precisely achieved using masks. By leveraging SAM (Kirillov et al., 2023) or other segmentation tools, users can generate a mask $m$ for important concepts within a given image $x$. The user-provided masks and original images are then collectively fed into the protected image generation module as described in Eq. (8). In this module, we propose a regional adversarial learning loss to reduce the visibility of protected images through precise protection and a Lagrangian multiplier-based solution to minimize perturbations while maintaining successful protection.
In this section, we start by formulating the regional adversarial learning framework for diffusion models in Sec. 4.2 and then move on to the solution of the proposed optimization objectives in Sec. 4.3. The total learning process is shown in Alg. 1.
Figure 2. The pipeline of VCPro. Initially, users can utilize SAM or alternative segmentation tools to create user-protective masks within images. Following this, the masks and the original images are simultaneously input into the protected image generation module, as described by Eq. (8). Within this module, we introduce a regional adversarial learning loss aimed at diminishing the visibility of protected images by precise protection and transform objectives to minimize perturbations while maintaining effective protection.
4.2. Formulation
Regional Adversarial Loss. Unlike previous studies, we aim to achieve precise concept protection to reduce the perceptibility of protected images. We use mask $m$ to indicate the spatial positions of important information in the feature map enabling precise concept protection. This precise optimization allows prioritized protection of the most critical information in the image with a smaller adversarial perturbation. The optimization objective of protected images in our method can be defined as:
(5) $\begin{gathered}x^{\prime}:=\arg\max_{x^{\prime}}\mathcal{L}^{\prime}_{\theta}\left(x^{\prime},y,m\right),\\ \text{ s.t. }\quad\left\|x-x^{\prime}\right\|\leq\epsilon,\\ \end{gathered}$
and the regional adversarial loss $\mathcal{L}^{\prime}_{\theta}$ is :
(6) $\mathcal{L}^{\prime}_{\theta}:=$ $\mathbb{E}_{z\sim\mathcal{E}(x),y,g\sim\mathcal{N}(0,1),t}\left[l_{mask}(z_{t},y,g,t,m)\right],$
$\text{where }\ l_{mask}(z_{t},y,g,t,m):=$ $\left\|g\odot m-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y,v)\right)\odot m\right\|_{2}^{2}-\left\|g\odot(1-m)-\varepsilon_{\theta}\left(z_{t},t,\tau_{\theta}(y,v)\right)\odot(1-m)\right\|_{2}^{2}.$
In that way, protected image $x^{\prime}$ pushes the output of the diffusion models far away from the ground truth Gaussian noise $g$ within a specified mask region, while predictions in other areas pull close. Without the pull component of the regional adversarial loss, the part outside the mask can not provide gradient feedback, making it difficult to optimize. It affects the optimization against perturbations, so we add the pull module, and removing the pull module will affect the protection effect as shown in Table 3.
4.3. Lagrangian Multiplier-based Loose Solution
Based on Eq. (6), protected images will destroy the target area as much as possible. However, for the generation task, the output can still be recognized as synthetic as long as there are some clear signs of protection. Typical image classification attacking methods (Carlini and Wagner, 2017; Luo et al., 2022; Chen et al., 2023) also present similar viewpoints that minimize the impact on normal visual perception while maintaining the effectiveness of the attack. Hence, we propose a loosed optimization objective by attempting to find the minimal adversarial perturbation $\delta$ that can attack diffusion models successfully. For convenience, we set a heuristic method: if $\mathcal{L}^{\prime}_{\theta}>\alpha$, the attack is successful. The problem can be translated as:
(7) minimize $D(x,x+\delta),$
such that $-\mathcal{L}^{\prime}_{\theta}+\alpha\leq 0,$
$\delta\in[-\epsilon,\epsilon]^{n},$
$x+\delta\in[0,255]^{n},$
where $D(\cdot)$ is a distance metric.
For ease of solution, we use Lagrangian multiplier method and get the alternative formulation:
(8) minimize $\mathcal{L}_{final}=D(x,x+\delta)+c\cdot(-\mathcal{L}^{\prime}_{\theta}+\alpha),$
such that $\delta\in[-\epsilon,\epsilon]^{n},$
$x+\delta\in[0,255]^{n},$
where a constant $c>0$ is appropriately selected. The equivalence of the Eq. (7) and Eq. (8) can be understood by the existence of a positive constant $c$ ensuring the best solution for the second formulation aligns with that of the first.
Considering that the human visual system is more sensitive to low-frequency regions (Luo et al., 2022), we use $D(\cdot)$ to limit perturbations into the high-frequency regions. Specifically, we use discrete wavelet transform (DWT) to transform images from the spatial domain to the frequency domain. DWT will decompose the image $x$ into one low-frequency and three high-frequency components, i.e., $x_{ll},x_{lh},x_{hl},x_{hh}$, and inverse DWT (IDWT) uses all four components to reconstruct the image.
$x_{ll}=LxL^{T},\quad x_{lh}=HxL^{T},$
$x_{hl}=LxH^{T},\quad x_{hh}=HxH^{T},$
where $L$ and $H$ are an orthogonal wavelet’s low-pass and high-pass filters, respectively. $x_{ll}$ preserves the low-frequency information of the original image, whereas $x_{lh},x_{hl}$ and $x_{hh}$ are associated with edges and drastic variations.
In this work, we drop the high-frequency components and reconstruct an image with only the low-frequency component as $\widetilde{x}=\phi(x)$, where $\phi(x)=L^{T}x_{ll}L=L^{T}(LxL^{T})L,\text{and }D(\cdot)=||\widetilde{x}-\widetilde{x^{\prime}}||_{2}^{2}$.
Algorithm 1 Visual-Friendly concept protection (VCPro) framework
0:  Image $x$, diffusion model with parameter $\theta$, number of time steps $T$, text condition $y$, attack steps $T_{a}$, step size $\alpha$, and adversarial perturbation size $\epsilon$
1:  Initialize $x^{\prime}=x$, $i=0$
2:  Get mask $m$ by SAM or user-provide
3:  while $i<T_{a}$ do
4:     Sample $t\sim[0,T]^{n}$
5:     $\delta=\operatorname{Uniform}(-\epsilon,\epsilon)$
6:     Calculate $\mathcal{L}_{final}(x^{\prime},y,m)$ by Eq. (8)
7:     $\delta=\delta-\alpha\cdot\operatorname{sign}\left(\nabla_{\delta}\mathcal{L}_{final}(x^{\prime},y,m)\right)$
8:     $\delta=\max(\min(\delta,\epsilon),-\epsilon)$
9:     $x^{\prime}=x^{\prime}+\delta$
10:     $x^{\prime}=\max(\min(x^{\prime},255),0)$
11:     $i=i+1$
12:  end while
13:  return Protected image $x^{\prime}$
5. Experiments
This section introduces our experimental settings and qualitative and quantitative experiments to demonstrate our effectiveness in generating protected images with less visual perceptibility while safeguarding key concepts in user-provided images.
5.1. Experimental Settings
Datasets. We used the CelebA-HQ (Karras et al., 2017) and VGGFace2 (Cao et al., 2018) datasets, following Anti-DreamBooth (Le et al., 2023) and MetaCloak (Liu et al., 2024), to validate the effectiveness of our method in privacy protection. CelebA-HQ is an enhanced version of the original CelebA dataset consisting of $30,000$ celebrity face images. VGGFace2 is a large-scale dataset with over $3.3$ million face images from $9,131$ unique identities. We randomly selected 50 identities from each dataset, ensuring that each had at least 15 images with a resolution exceeding 500 × 500 pixels. We also provide other non-face cases (artwork and landscape) in the paper.
Training details. The availability of Stable Diffusion pre-trained weights on Hugging Face  [^1] https://huggingface.co/ has significantly advanced research within the community. Here, our experiments primarily focus on the Stable Diffusion v1-4 version. For image pre-processing, we center crop and resize images to a resolution $512\times 512$. During protected image generation, we use $720$ iterations, step size $\alpha$ is $1/255$, and adversarial perturbation size $\epsilon$ is $8/255$. We test the protection performance on two types of SOTA personalization methods, i.e., Textual Inversion (TI) (Gal et al., 2022), and DreamBooth (DB) (Ruiz et al., 2023). During the training of the Textual Inversion, the constant learning rate is set to $5\times 10^{-4}$ with $3000$ optimization steps and batch size $1$. In DreamBooth, the constant learning rate is set to $5\times 10^{-7}$ with $1000$ optimization steps and batch size is $2$. We use a guidance scale of $7.5$ and $50$ denoising steps at the generation phase with a testing prompt. We set $c=0.1$ and $\alpha=0.5$ in practice. All experiments are complete on NVIDIA A100 GPU 40GB.
Baselines. We compare our method with five SOTA baselines, i.e., Mist (Liang et al., 2023), Anti-DreamBooth (referred to as Anti-DB) (Le et al., 2023), PhotoGuard (Salman et al., 2023), SDS(-) (Xue et al., 2024) and MetaCloak (Liu et al., 2024). In implementations, we set $\epsilon=8/255$ as the $\ell_{\infty}$ perturbation size, $\alpha=1/255$ as the step size for fair comparison.
  • Mist aims to maximize the training loss of Latent Diffusion Models (LDMs) while minimizing a textual loss between the protected image and a dummy image in the image encoder layer during protected image generation. In our implementation, following the original paper (Liang and Wu, 2023), we set the loss weight for the textual loss to $1\times 10^{-4}$ and run 720 iterations during protected image generation.
  • Anti-DreamBooth focuses on maximizing the training loss of LDMs to generate protected images. We use the most visually imperceptible variant, FSMG, which employs a DreamBooth model trained on original images for protected data generation. The adversarial iterations were set to 720.
  • PhotoGuard targets diffusion-based inpainting/editing methods and offers two variants: one that attacks the VAE-encoder of LDMs, and another that attacks the final image outputs. We select the VAE-encoder variant due to its superior visual quality and set the training iterations to 200, as recommended in the original paper (Salman et al., 2023).
  • SDS(-) minimizes the SDS loss (Poole et al., 2022) to optimize adversarial data, achieving the best visual quality among all versions presented in its original paper (Xue et al., 2024). We set the training iterations to 100, following (Xue et al., 2024).
  • MetaCloak utilizes meta-learning to enhance the robustness of adversarial data against various transformations. We set the number of surrogate models to 5, the unrolling number to 1, the sample batch size to 1, and the training iterations to 4000, based on the recommendations in (Liu et al., 2024).
[…]
Table 1. Automatic and human evaluations across different methods, including their respective mean and standard deviation ($\pm$) on the VGGFace2 and CelebA-HQ datasets. All methods employed a perturbation size of $\epsilon$ = 8/255. In the HumanEval results, “Loss/Tie/Win” in “Visibility” indicates the voting rates for VCPro losses, ties, and wins in comparison to other methods regarding the stealthiness of their adversarial perturbations. “TI/DB” represents the percentage of images generated by Textual Inversion/DreamBooth identified as synthetic. ↑ means the higher the better, and vice versa. The best result in each column is in bold.
Methods Publication Visibility of Perturbations Textual Inversion DreamBooth HumanEval
FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$ Visibility (Loss/Tie/Win) TI $\uparrow$ DB $\uparrow$
VGGFace2 No Defense - - - - $3.61_{\pm 0.98}$ $0.24_{\pm 0.34}$ $3.95_{\pm 1.11}$ $0.38_{\pm 0.25}$ 64 : 26 : 10 - -
Mist arXiv23 $105.93_{\pm 9.76\hphantom{0}}$ $0.83_{\pm 0.03}$ $32.48_{\pm 1.12}$ $1.24_{\pm 0.41}$ $0.11_{\pm 0.29}$ $1.02_{\pm 0.11}$ $0.25_{\pm 0.28}$ 04 : 08 : 88 098.65 097.26
Anti-DreamBooth ICCV23 $\hphantom{0}96.24_{\pm 10.44}$ $0.82_{\pm 0.03}$ $32.43_{\pm 1.10}$ $1.18_{\pm 0.36}$ $0.05_{\pm 0.29}$ $1.03_{\pm 0.14}$ $0.26_{\pm 0.28}$ 03 : 09 : 88 097.33 100.00
PhotoGuard ICML23 $\hphantom{0}62.07_{\pm 8.15\hphantom{0}}$ $0.84_{\pm 0.03}$ $33.17_{\pm 1.21}$ $1.50_{\pm 0.54}$ $0.13_{\pm 0.34}$ $1.26_{\pm 0.30}$ $0.27_{\pm 0.31}$ 16 : 21 : 63 098.72 083.95
SDS(-) ICLR24 $\hphantom{0}47.48_{\pm 6.81\hphantom{0}}$ $0.81_{\pm 0.03}$ $32.91_{\pm 0.43}$ $2.09_{\pm 0.82}$ $0.07_{\pm 0.37}$ $2.74_{\pm 0.97}$ $0.33_{\pm 0.26}$ 14 : 17 : 69 094.44 077.03
MetaCloak CVPR24 $204.25_{\pm 54.52}$ $0.82_{\pm 0.02}$ $32.09_{\pm 0.63}$ $1.86_{\pm 0.62}$ $0.19_{\pm 0.33}$ $1.33_{\pm 0.41}$ $0.25_{\pm 0.32}$ 02 : 13 : 85 091.18 090.41
VCPro $\hphantom{0}27.04_{\pm 4.38\hphantom{0}}$ $0.90_{\pm 0.03}$ $35.33_{\pm 1.79}$ $2.31_{\pm 0.78}$ $0.03_{\pm 0.32}$ $2.06_{\pm 0.70}$ $0.21_{\pm 0.34}$ 098.57 097.37
CelebA-HQ No Defense - - - $4.40_{\pm 0.86}$ $0.41_{\pm 0.22}$ $4.84_{\pm 0.40}$ $0.63_{\pm 0.13}$ 79 : 18 : 03 - -
Mist arXiv23 $\hphantom{0}78.32_{\pm 11.83}$ $0.86_{\pm 0.03}$ $33.75_{\pm 0.21}$ $1.45_{\pm 0.58}$ $0.26_{\pm 0.26}$ $1.04_{\pm 0.11}$ $0.44_{\pm 0.21}$ 06 : 03 : 91 096.20 100.00
Anti-DreamBooth ICCV23 $\hphantom{0}78.45_{\pm 11.66}$ $0.86_{\pm 0.03}$ $33.74_{\pm 0.21}$ $1.81_{\pm 1.20}$ $0.28_{\pm 0.25}$ $1.04_{\pm 0.15}$ $0.44_{\pm 0.21}$ 00 : 05 : 95 096.92 096.51
PhotoGuard ICML23 $\hphantom{0}45.18_{\pm 7.21\hphantom{0}}$ $0.88_{\pm 0.02}$ $35.08_{\pm 0.25}$ $1.81_{\pm 0.59}$ $0.30_{\pm 0.24}$ $1.23_{\pm 0.35}$ $0.48_{\pm 0.17}$ 04 : 21 : 75 094.12 094.37
SDS(-) ICLR24 $\hphantom{0}34.53_{\pm 6.16\hphantom{0}}$ $0.82_{\pm 0.03}$ $33.01_{\pm 0.28}$ $2.29_{\pm 0.70}$ $0.09_{\pm 0.28}$ $2.92_{\pm 0.74}$ $0.55_{\pm 0.15}$ 09 : 10 : 81 098.70 084.51
MetaCloak CVPR24 $161.93_{\pm 26.93}$ $0.86_{\pm 0.03}$ $33.12_{\pm 0.20}$ $1.47_{\pm 0.56}$ $0.25_{\pm 0.26}$ $1.59_{\pm 0.53}$ $0.44_{\pm 0.16}$ 00 : 05 : 95 096.00 095.52
VCPro $\hphantom{0}16.21_{\pm 2.89\hphantom{0}}$ $0.95_{\pm 0.01}$ $39.32_{\pm 0.40}$ $2.61_{\pm 1.00}$ $0.09_{\pm 0.28}$ $2.62_{\pm 0.95}$ $0.44_{\pm 0.19}$ 100.00 095.64
As image quality assessment remains a challenging problem, particularly with the lack of precise concept-specific automated metrics, we also provide human evaluation to assess the visual perception of protected data and the efficacy of safeguarding the target concepts.
Figure 3. Qualitative defense results for both the human face and non-face cases. For generating protected images, we use an $\epsilon$ value of 8/255 with Stable Diffusion v1-4. Each row displays the original images (No Defense), the protected images generated by PhotoGuard, Anti-DreamBooth, Mist, SDS(-), MetaCloak, and our method, along with their respective outcomes in Textual Inversion (TI) and DreamBooth (DB). Additionally, we include visualizations of adversarial perturbations represented as black-yellow images. For better detail, please zoom in.
5.2. Main Results
Quantitative Results.
Table 1 shows the automatic and human evaluation results for the visibility of protected images and the final protection effectiveness of different methods.
For visibility of protected images, we evaluate the average FID/SSIM/PSNR differences between all protected images and their corresponding original images. The statistics show that the protected images generated by our method are more faithful to the original images, and the least variance of FID shows that we also have a stable effect. Compared with our baseline Anti-DreamBooth, we can max reduce 69.20 in FID, up to 0.09 in SSIM, and up to 5.58 in PSNR. Compared with current SOTA methods PhotoGuard and SDS (-), our image quality is still significantly improved. Although SDS (-) achieves a low FID, it introduces noticeable brightness increases in the perturbations, leading to poor SSIM performance.
Protected images against unauthorized AI generation by improving training costs and reducing final AI-generated image quality. For final protection effectiveness, we still have significant relative decreases in full-image quality metric LIQE compared to the no defense (from “good/perfect” to “poor”), again achieving a protective effect, which is similarly confirmed by the manual evaluation later on human evaluation and qualitative results. Considering that the full-image quality metric is not fairly for us, we also add a face region quality metric CLIP-FACE. In CLIP-FACE, we achieve the best protection in most cases. Both metrics demonstrate that we have reduced visibility against perturbations while still having a strong protection effect.
Qualititative Results. Fig. 3 illustrates protected images and protective effects on individual identity information and landscapes of different methods under Textual Inversion and DreamBooth. We also visualize adversarial perturbations by regularizing all perturbations to the range [0,1] and then visualizing using colormap.
Under the protection of our method, the personalization methods fail to generate visual-fidelity results. As shown in the case of important information such as faces, landmark landscapes, or buildings, our adversarial perturbation is less obvious (especially in the background part), and effectively distorts textures and features in important regions, thereby preventing the malicious personalization of individual identification information or copyright, so that the adversarial perturbation is more subtle while achieving the protection purpose, and find a better balance between the protection effect and the perturbation intensity, which reduces user concerns about image quality degradation after protection. PhotoGuard, Mist, Anti-DreamBooth, and MetaCloak tend to add strange textures to the whole image and show obvious adversarial perturbations (the second, third, fourth, and sixth columns). SDS(-) can cause generated images to be too blurry and has limited protection for the DreamBooth. Its protected image is like adding strange circular blobs to the original image (the fifth column). Visualization of its perturbations also supports the idea that excessive brightness leads to lower SSIM values. These approaches serve as a protective measure, ensuring that certain image regions exhibit unique textures. The presence of such textures can strongly suggest infringement. We also find that the protection difficulty of Textual Inversion is far less than that of DreamBooth because DreamBooth usually fine-tunes most parameters in Stable Diffusion. At the same time, Textual Inversion only adds a few new word embeddings. It is also supported in Table 1.
User Study. We evaluate the performance of VCPro in the concept protection task with five selected methods, i.e., PhotoGuard (Salman et al., 2023), Mist (Liang and Wu, 2023), Anti-DreamBooth (Le et al., 2023), SDS(-)(Xue et al., 2024), and MetaCloak (Liu et al., 2024). A total of 27 participants took part in the survey, and they all use social media proficiently. We randomly sampled protected images and generation images from all 100 identities in VGGFace2 and CelebA-HQ and six groups of non-face data. The study consists of three parts, including visibility of adversarial perturbations (1,116 valid votes), protection effects under Textual Inversion (1,078 valid votes), and protection effects under DreamBooth (1,067 valid votes).
User Study I
In the survey on the visibility of adversarial perturbations, participants were shown a VCPro-protected image alongside a protected image generated by another method based on the same original image. They were then asked to vote on the visual quality, choosing between ”A wins,” ”tie,” or ”B wins.” The visual quality of VCPro was significantly better than that of other methods. Specifically, when compared to Mist, VCPro received 88.27% of the votes (6.79% ties and 4.94% for Mist); against Anti-DreamBooth, VCPro received 91.72% (7.01% ties and 1.27% for Anti-DreamBooth); against MetaCloak, VCPro received 90.68% (8.70% ties and 0.62% for MetaCloak); against PhotoGuard, VCPro received 70.25% (20.89% ties and 8.86% for PhotoGuard); and against SDS(-), VCPro received 76.97% (12.73% ties and 10.30% for SDS(-)).
User Study II
In the survey on protection efficacy under Textual Inversion, participants were given an original image $x$ and asked to determine whether the image generated through Textual Inversion, trained on protected image $x^{\prime}$ from different methods, was synthetic. The results were as follows: for Mist, 98.58% voted synthetic; for Anti-DreamBooth, 96.08%; for MetaCloak, 94.19%; for PhotoGuard, 96.79%; for SDS(-), 96.08%; and for VCPro, 98.66%.
User Study III
In the survey on protection efficacy under DreamBooth, participants were shown an original image and a DreamBooth-generated image trained on its protected images from different methods. They were then asked to determine whether the latter was synthetic. The results were as follows: for Mist, 98.50% voted synthetic; for Anti-DreamBooth, 98.00%; for MetaCloak, 92.86%; for PhotoGuard, 86.93%; for SDS(-), 83.55%; and for VCPro, 95.21%. Most participants also acknowledged that VCPro had a protective effect.
[…]
5.3. Perturbation Size Influence
Perturbation size $\epsilon$ controls the maximum allowable change in pixel values of adversarial perturbation. As shown in Fig. 4 and Table 2, different levels of adversarial perturbation size noticeably influence the protection outcomes: Compared to low $\epsilon$, large $\epsilon$ presents worse invisibility of adversarial perturbations while the more obvious protection effects.
For Textual Inversion, when the perturbation size is low, the image preserves the facial area with alterations in facial texture and feature distribution. Upon reaching a perturbation size of 16/255, the facial areas experience complete degradation. The results under DreamBooth and quantitative experiment show a similar trend. Compared with DreamBooth, Textual Inversion is easier to achieve concept protection.
We visualize adversarial perturbations by normalizing all perturbations to the range [0,1] and applying a colormap. An interesting pattern emerges: as $\epsilon$ increases, perturbations spread from the face to the background. This diffusion occurs because, when perturbation size is constrained, the perturbations are concentrated in areas that most effectively enhance the protective effect.
Table 2. The influence of different perturbation size $\epsilon$ on VGGFace2 datasets. ↑ means the higher the better, and vice versa. The best result in each column is in bold.
Budget Visibility of Perturbations Textual Inversion DreamBooth
FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$
0 - - - 3.61±0.98 00.24±0.34 3.95±1.11 0.38±0.25
4 22.98±4.08 0.93±0.03 36.45±2.21 2.76±0.77 -0.02±0.27 2.33±0.70 0.25±0.31
8 27.04±4.38 0.90±0.03 35.33±1.79 2.31±0.78 00.03±0.32 2.06±0.70 0.21±0.34
16 48.67±9.02 0.82±0.03 32.63±1.06 1.86±0.80 -0.02±0.23 1.81±0.53 0.19±0.32
Figure 5. The ablation experiments of different mask sizes and modules. Anti-DreamBooth (Anti-DB) is our baseline. The red region indicates the important region. Here Input-mask means optimizing adversarial perturbations by directly restricting the mask range. Zoom in for a better view.
5.4. Ablation Study
As shown in Fig. 5 and Table 3, we conducted ablation experiments of two modules in VCPro: Regional Adversarial Loss (RAL) in Sec. 4.2 and Lagrangian Multiplier-based Loose Solution (LMLS) in Sec. 4.3.
Compared with our baseline Anti-DreamBooth, we find that Anti-DreamBooth+RAL reduces the visibility of the perturbations, while the protection effect is concentrated in the mask area, resulting in a slight decrease in the global image quality assessment LIQE but stable in CLIP-FACE. At the same time, we also consider the effect of different mask ranges, smaller masks will further reduce the visibility of the perturbations, and the protection range is further reduced.
LMLS shows a similar phenomenon, where Anti-DreamBooth+LMLS reduces the visibility of the noise and tends to leave the high-frequency part, while the final protection effect also leaves a streaky texture across the whole image. That is to say, the adversarial perturbation added to a specific space and frequency domain will directly affect the protection effect in the corresponding space and frequency domains.
Moreover, only push-loss without pull-loss in RAL (VCPro (w/o Pull) in Fig. 5 and Table 3) will reduce the training loss gradient feedback about the protection effect, while the loss that constrains the visibility of the perturbations is not affected resulting in decreased visibility of the perturbations and the protection effect.
We also compare the difference between directly protected images updated within the mask range (Input-Mask) and us adding guidance at the loss side, both can achieve the protection of the specified region. When adding the mask constraint directly at the protected images, it improves pixel-level global metrics such as SSIM/PSNR. However, it makes the noise of the protected target on the original image more obvious than us and FID is still high. We can achieve a better visual perception by combining the two, and at the same time, achieve the protection of the target.
Table 3. Automatic evaluation results on ablation experiments with the corresponding mean and standard deviation ($\pm$). All methods use $\epsilon$ = 8/255 on VGGFace2 dataset. ↑ means the higher the better, and vice versa. The best result in each column is in bold. Here Input-Mask is to constrain the adversarial perturbations directly on the input image to be only within the mask, we are constraining on the optimization objective, and the adversarial perturbations will be present in the whole image.
Methods Visibility of Perturbations Textual Inversion DreamBooth
FID $\downarrow$ SSIM $\uparrow$ PSNR $\uparrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$
No Defense - - - $3.61_{\pm 0.98}$ $0.24_{\pm 0.34}$ $3.95_{\pm 1.11}$ $0.38_{\pm 0.25}$
Anti-DreamBooth $96.24_{\pm 10.44}$ $0.82_{\pm 0.03}$ $32.43_{\pm 1.10}$ $1.18_{\pm 0.36}$ $0.05_{\pm 0.29}$ $1.03_{\pm 0.14}$ $0.26_{\pm 0.28}$
Anti-DreamBooth+RAL $68.18_{\pm 8.58\hphantom{0}}$ $0.83_{\pm 0.03}$ $32.89_{\pm 1.25}$ $1.62_{\pm 0.57}$ $0.03_{\pm 0.31}$ $1.53_{\pm 0.53}$ $0.18_{\pm 0.31}$
Anti-DreamBooth+RAL (Small Mask) $58.44_{\pm 9.34\hphantom{0}}$ $0.84_{\pm 0.03}$ $33.09_{\pm 1.27}$ $2.11_{\pm 0.82}$ $0.05_{\pm 0.33}$ $2.21_{\pm 0.66}$ $0.27_{\pm 0.32}$
Anti-DreamBooth+LMLS $30.08_{\pm 5.22\hphantom{0}}$ $0.90_{\pm 0.03}$ $35.03_{\pm 1.83}$ $2.52_{\pm 0.82}$ $0.17_{\pm 0.34}$ $2.16_{\pm 0.65}$ $0.30_{\pm 0.30}$
VCPro (w/o Pull) $17.21_{\pm 3.13\hphantom{0}}$ $0.94_{\pm 0.03}$ $36.79_{\pm 2.34}$ $2.74_{\pm 0.82}$ $0.14_{\pm 0.33}$ $2.55_{\pm 0.92}$ $0.35_{\pm 0.32}$
VCPro $27.04_{\pm 4.38\hphantom{0}}$ $0.90_{\pm 0.03}$ $35.33_{\pm 1.79}$ $2.31_{\pm 0.78}$ $0.03_{\pm 0.32}$ $2.06_{\pm 0.70}$ $0.21_{\pm 0.34}$
Anti-DreamBooth+Input-Mask $41.27_{\pm 8.76\hphantom{0}}$ $0.90_{\pm 0.03}$ $35.15_{\pm 2.02}$ $2.05_{\pm 0.76}$ $0.10_{\pm 0.32}$ $2.03_{\pm 0.80}$ $0.23_{\pm 0.33}$
VCPro+Input-Mask $18.61_{\pm 3.19\hphantom{0}}$ $0.93_{\pm 0.03}$ $36.61_{\pm 2.25}$ $2.41_{\pm 0.78}$ $0.07_{\pm 0.34}$ $2.68_{\pm 1.05}$ $0.31_{\pm 0.32}$
Sensitivity Analysis of Training Interation
Usually the larger the training steps, the larger the value of $\mathcal{L}^{\prime}_{\theta}$ at different time steps, and the more successful the protection is. As shown in Fig. LABEL:fig:hyp (a), when increasing the number of training iterations, FID will first improve rapidly and then rise slowly, and the overall trend is that the longer the training iterations, the worse the quality of the protected image. Whereas LIQE will fall rapidly and then level off, the protection effect will improve and then keep stable.
Sensitivity Analysis of $c$
As shown in Fig. LABEL:fig:hyp (b), as the parameter $c$ increases, representing a stronger constraint on perturbation visibility, the FID score decreases, indicating improved visibility of the final protected image. Correspondingly, the image protection effectiveness (LIQE) initially decreases, then fluctuates around 5.0. Based on the balance between protection effectiveness and noise quality, we ultimately select $c=0.1$.
5.5. Analysis of Adversary Settings
Following Anti-DreamBooth (Le et al., 2023), we consider three settings in this paper: “convenient setting”, “adverse setting”, and “uncontrolled setting”.
Figure 6. Experiments about different adversarial settings. Zoom in for a better view.
Convenient Setting. In a convenient setting (white-box setting), we (protectors) have all knowledge of the resources of adversary, including a pre-trained text-to-image generator, training term (e.g., “sks”), and training prompt. This scenario is practical because adversaries are likely to use high-quality, open-source pre-trained generators, with Stable Diffusion being the most prominent publicly available model. Additionally, people often rely on open-source code with default training terms and prompts. The convenient setting is the simplest for the protectors, and all relevant experiment results have been analyzed and mentioned in Sec. 5.2.
Adverse Settings. In adverse settings, also known as grey-box settings, protectors lack knowledge of the specific version of the pre-trained text-to-image generator, training term, or training prompt employed by the adversary. This scenario more closely resembles real-world conditions. To defend against such attacks, protectors typically use surrogate modules to test the transferability of adversarial perturbations. Experiments under both assumptions demonstrate the practicality of our method in real-world scenarios. All experiments utilize $\epsilon$ = 8/255 as the perturbation size on the VGGFace2 dataset, with Stable Diffusion v1-4 as the base model, except in the case of model mismatching.
  • Model Mismatching. In this scenario, protectors use Stable Diffusion v1-4 without knowing the specific version of Stable Diffusion employed by the adversary. Results in Fig. 6 and Table 4 show effective facial area protection even if adversaries used Stable Diffusion v1-4/v1-5/v2-1, which reflects the transferability of VCPro.
  • Training Term Mismatching. We set protectors to use “sks” as a training term in Textual Inversion and DreamBooth, and adversaries use “t@t”. As shown in Fig. 6 and Table 5, VCPro is robust to different training terms.
  • Training Prompt Mismatching. We set protectors to use “a photo of sks person.” as training prompt in Textual Inversion and DreamBooth, and adversaries to use “a dslr portrait of sks person.”. As shown in Fig. 6 and Table 6, VCPro can still successfully protect the important concepts.
  • Post-Processing. In this setting, we assume that the adversaries will use some post-processing to reduce the impact of the protected images, taking the two common post-processing techniques, Gaussian filtering and JPEG compression, as an example. Here we set the kernel size $7\times 7$ for Gaussian filtering and adopt 75% quality compression for JPEG, which also mentioned in (Sandoval-Segura et al., 2023) can be a good purification method. As shown in Fig. 6 and Table 7, although the protective effect is somewhat diminished, the final generated image is still visibly cracked. VCPro can still successfully protect the target area.
Uncontrolled Settings. In this section, we examine the scenario where an adversary has access to a set of original images of the target concept, which are then combined with protected images to train DreamBooth. We evaluate three settings in which the proportion of original images is varied from 25% to 75%, as detailed in Table 8. All experiments employ a perturbation size of $\epsilon$ = 8/255 on the VGGFace2 dataset and utilize Stable Diffusion v1-4 as the base model. Our protection remains effective as long as the proportion of protected images exceeds 50%, but its effectiveness diminishes as the proportion of original images increases. Ideally, VCPro would receive government support, become widely adopted, and be implemented across all social media platforms, thereby mitigating the risks associated with these uncontrolled settings.
Table 4. Results of model mismatching setting with the corresponding mean and standard deviation ($\pm$). ↑ denotes that higher values are preferable, while ↓ indicates the opposite. The best result in each column is highlighted in bold.
Protector Adversary Textual Inversion DreamBooth
LIQE $\downarrow$ CLIP-FACE $\downarrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$
No Defense Stable Diffusion v1-4 3.61±0.98 0.24±0.34 3.95±1.11 0.38±0.25
Stable Diffusion v1-4 Stable Diffusion v1-4 2.31±0.78 0.03±0.32 2.06±0.70 0.21±0.34
Stable Diffusion v1-4 Stable Diffusion v1-5 2.06±0.74 0.02±0.32 2.00±0.57 0.24±0.34
Stable Diffusion v1-4 Stable Diffusion v2-1 base 1.82±0.60 0.06±0.30 1.70±0.49 0.20±0.34
Table 5. Results of training term mismatching setting with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
Protector Adversary Textual Inversion DreamBooth
LIQE $\downarrow$ CLIP-FACE $\downarrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$
No Defense sks 3.61±0.98 0.24±0.34 3.95±1.11 0.38±0.25
sks sks 2.31±0.78 0.03±0.32 2.06±0.70 0.21±0.34
sks t@t 2.03±0.74 0.04±0.34 1.93±0.66 0.21±0.33
Table 6. Results of training prompt mismatching setting with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
Protector Adversary Textual Inversion DreamBooth
LIQE $\downarrow$ CLIP-FACE $\downarrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$
No Defense “a photo of sks person.” 3.61±0.98 0.24±0.34 3.95±1.11 0.38±0.25
“a photo of sks person.” “a photo of sks person.” 2.31±0.78 0.03±0.32 2.06±0.70 0.21±0.34
“a photo of sks person.” “a dslr portrait of sks person.” 2.23±0.77 0.04±0.32 2.62±1.18 0.22±0.32
Table 7. Results of different post-processing with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
Protector Adversary Textual Inversion DreamBooth
LIQE $\downarrow$ CLIP-FACE $\downarrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$
No Defense Null 3.61±0.98 0.24±0.34 3.95±1.11 0.38±0.25
Null Null 2.31±0.78 0.03±0.32 2.06±0.70 0.21±0.34
Null Gaussian Filtering 2.56±0.78 0.18±0.31 2.18±0.44 0.31±0.34
Null JPEG Compression 2.08±0.79 0.12±0.33 2.09±0.46 0.29±0.32
Table 8. Results of different uncontrolled settings with the corresponding mean and standard deviation ($\pm$). ↑ means the higher the better, and vice versa. The best result in each column is in bold.
Perturbed Clean Textual Inversion DreamBooth
LIQE $\downarrow$ CLIP-FACE $\downarrow$ LIQE $\downarrow$ CLIP-FACE $\downarrow$
0 100% 3.61±0.98 0.24±0.34 3.95±1.11 0.38±0.25
25% 75% 2.82±1.12 0.20±0.34 3.50±1.25 0.40±0.28
50% 50% 2.15±0.81 0.05±0.35 2.89±1.21 0.35±0.32
75% 25% 2.03±0.74 0.04±0.34 2.71±1.07 0.33±0.32
100% 0% 2.31±0.78 0.03±0.32 2.06±0.70 0.21±0.34
6. Conclusion
In this paper, we show that existing approaches utilizing adversarial perturbations to safeguard images from malicious personalization often overemphasize the final protection effectiveness, resulting in more noticeable perturbations. To mitigate this problem, on the one hand, we protect the important concept regions rather than the full images in previous works, leveraging the sparse nature of images and designing a user-specified image concepts protection framework. On the other hand, we change the optimization objective from generating the most protective adversarial perturbation to generating the least perceptible adversarial perturbation that exactly achieves the required protective effect. Quantitative and qualitative experiments demonstrate that we can protect important user-specified concepts and greatly reduce the degree of naked-eye visibility of adversarial perturbations.
Future Works and Limitations
We have not yet reached the capability to generate protected images in real time. However, we are committed to ongoing research and development in this area. Our future efforts will focus on finding efficient methods to produce protected images swiftly while maintaining high visual quality. By doing so, we aim to significantly enhance the overall user experience and ensure that our solutions meet the highest standards of both functionality and aesthetics.